OpenStudy (anonymous):

a ball is thrown vertically down with initial velocity 10 m/s from 15 m above the ground. calculate the velocity when it hits the ground!

OpenStudy (abhisar):

Hello @jansd345 , Welcome to OpenStudy !

OpenStudy (abhisar):

Here, initial velocity = 10 m/s Acceleration = 9.8 m/s Distance covered = 15 m We have to find final velocity. We can use the equation \(\sf v^2=u^2+2as\) \(\Rightarrow v^2=10^2+2\times 9.8 \times 15\) => \(\sf v^2\) = 394 => V= \(\sqrt{394}\) = 19.84 m/s

OpenStudy (anonymous):

thank you so much!

OpenStudy (abhisar):

Do you know you can reward those who help u by hitting the blue best response button. You're most welcome :)

OpenStudy (anonymous):

although I changed the 9.8 m/s to 10 m/s because thats what it says here g=10 m/s :) sure I'll click that button

OpenStudy (abhisar):

Yes u can change it to 10 m/s if they say so. There will be a slight difference in the final answer then :)

OpenStudy (abhisar):

It will be 20 m/s i think

OpenStudy (anonymous):

Yes you are right, I have another problem. A ball is thrown vertically downwards with initial velocity 5 m.s from 30 above the ground. how long does it take for the ball to hit the ground? g=10 m/s please and thank you :)

OpenStudy (anonymous):

and also a ball is dropped from a tower. when it lands, its velocity is 30 m/s g=10 m/s^2. how long does it take for the ball to hit the ground?

OpenStudy (abhisar):

ok you need to ask these questions in a new post.

OpenStudy (abhisar):

For that you will have to close this question first.